Match the following aqueous solutions with the appropriate letter from the column on the right. 1. 0.10 m AlCl3 A. Highest boiling point 2. 8.6×10-2 m Cr2(SO4)3 B. Second highest boiling point 3. 0.19 m NaNO3 C. Third highest boiling point 4. 0.46 m Glucose(nonelectrolyte) D. Lowest boiling point
Answer #1

1. 0.10 m AlCl3.....................> C. Third highest boiling point

AlCl3 .............> Al^3+ + 3 Cl-

molality of particle concentration = 4 * 0.10 = 0.40 m

2. 8.6×10-2 m Cr2(SO4)3 .....................> B. Second highest boiling point

Cr2(SO4)3................> 2 Cr^3+ + 3 SO4^2-

molality of particle concentration = 5 * 8.6×10-2 = 0.43 m

3. 0.19 m NaNO3 ..................> D. Lowest boiling point

NaNO3 ...........> Na+ + NO3-

molality of particle concentration = 0.19 * 2 = 0.38 m

4. 0.46 m Glucose (nonelectrolyte)..................> A. Highest boiling point

molality of particle concentration = 0.46 m